OpenStudy (anonymous):

*MEDAL* Show your work to calculate the discriminant of the quadratic equation, and then state the number and type of solutions you expect the equation to have, based on the value of your discriminant. 0=x^2+6x-16

OpenStudy (anonymous):

i am pretty sure you can compute \[b^2-4ac\] with \[a=1,b=6,c=-16\] i.e. compute \[6^2-4\times 1\times (-16)\]

OpenStudy (anonymous):

you could even use a calculator if you liked, but the next step would be \[36+64\]

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

I think so, with your help and a help video I'm watching I think I got it. Thank you so much!

OpenStudy (anonymous):

yw btw since 100 is a perfect square (it is the square of 10) that means your two solutions will be rational numbers integers in fact

OpenStudy (anonymous):

which also means you can factor \[ 0=x^2+6x-16\\ 0=(x+8)(x-2)\] etc

OpenStudy (anonymous):

Ohh alright I see

OpenStudy (anonymous):

if the discriminant was not a perfect square, say it was 10 or 15, then the two zeros would not be rational numbers and you would have to solve by some other method, like using the quadratic formula

OpenStudy (anonymous):

okay I worked it out on my own and got this. (6)^2-4(1)(-16) 36-4(-16) -4x-16=64 4+64=68 I think I messed up somewhere but I'm not sure where.

OpenStudy (anonymous):

@satellite73